OpenStudy (anonymous):

What is the perimeter of a triangle with the given side lengths? s1 = x + 4 cm s2 = 3x + 1 cm s3 = 7x + 3 cm

OpenStudy (asnaseer):

Do you know what "perimeter" means?

OpenStudy (anonymous):

of course lol

OpenStudy (asnaseer):

so where are you stuck?

OpenStudy (anonymous):

you need to add all the given side together since u are solving for perimeter

OpenStudy (anonymous):

s1+s2+s3

OpenStudy (anonymous):

with the variables in the equation.. I dont understand why they are there and what they are doing

OpenStudy (anonymous):

adding all the sides together you can find x which u would need to then substitute into each equation to find the perimeter

OpenStudy (asnaseer):

just add "like" items e.g \[(x+2)+(2x-1)=x+2+2x-1=(x+2x)+(2-1)=3x+1\]

OpenStudy (anonymous):

you just need to add the tree sides together and simplify

OpenStudy (anonymous):

so x+4+3x+1+7x+3 collect the like terms x+3x+7x+4+1+3 then solve

OpenStudy (asnaseer):

@linzee1877 - do you understand?

OpenStudy (anonymous):

yes exactly

OpenStudy (anonymous):

yeah I think so

OpenStudy (anonymous):

thank you all

OpenStudy (asnaseer):

think of it like adding different "types" of fuit together. imagine the "x"'s were apples and the "constants" (i.e. the numbers on their own) were oranges. if we said x+1+2x+3 you could think of it as 1apple plus 1 orange + 2 apples plus 3 oranges so you have 3 apples and 4 oranges altogether (i.e. 3x + 4)

OpenStudy (asnaseer):

hope that helps :)

OpenStudy (anonymous):

it does, thank you

OpenStudy (asnaseer):

yw :)